given: f(x)=x^2+3 and g(x) = (sqrt)2-x find f o g(x)
Answer by Nate(3500)

f(g(x)) means that the g(x) is placed in the f(x) function
f(sqrt(2 - x)) = (sqrt(2 - x))^2 + 3
f(g(x)) = 2 - x + 3 = -x + 5